# First is Not Really Better Than Last: Evaluating Layer Choice and Aggregation Strategies in Language Model Data Influence Estimation

### v1

@@ Abstract
Identifying how training samples influence/impact Large Language Model (LLM) decision-making is essential for effectively interpreting model decisions and auditing large-scale datasets. Current training sample influence estimation methods (also known as influence functions) undertake this goal by utilizing information flow through the model via its first-order and higher-order gradient terms. However, owing to the large model sizes of today consisting of billions of parameters, these influence computations are often restricted to some subset of model layers to ensure computational feasibility. Prior seminal work by Yen-2022 in assessing which layers are best suited for computing language data influence concluded that the first (embedding) layers are the most informative for this purpose, using a hypothesis based on influence scores canceling out (i.e., the cancellation effect). In this work, we propose theoretical and empirical evidence demonstrating how the cancellation effect is unreliable, and that middle attention layers are better estimators for influence. Furthermore, we address the broader challenge of aggregating influence scores across layers, and showcase how alternatives to standard averaging (such as ranking and vote-based methods) can lead to significantly improved performance. Finally, we propose better methods for evaluating influence score efficacy in LLMs without undertaking model retraining, and propose a new metric known as the Noise Detection Rate (NDR) that exhibits strong predictive capability compared to the cancellation effect. Through extensive experiments across LLMs of varying types and scales, we concretely determine that the first (layers) are not necessarily better than the last (layers) for LLM influence estimation, contrasting with prior knowledge in the field.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated stellar performance on tasks across a number of applications and domains (Street-2024; Mittelst-2024; Marco-2025). Despite these advancements, current state-of-the-art models still exhibit suboptimal behavior on complex reasoning tasks (Jiang-2025), hallucinate responses and facts (Cleti-2024), and can make biased and unfair decisions (Gallegos-2024; Peters-2025). To improve the trust and safety of LLMs, it is imperative to better interpret and understand model decision-making (Singh-2024).
Recently, data attribution and influence methods (Hammoudeh-2024) have shown great promise in interpreting LLM behavior from the perspective of the training data (Grosse-2023; Chhabra-2025). These approaches seek to detect noisy, anomalous, out-of-distribution, or problematic training samples and conceptually link them to model output performance (Pleiss-2020; Yang-2024; Jiang-2021). Even highly curated datasets can contain detrimental samples, often introduced unknowingly through human error or surreptitiously by malicious adversaries (Ekambaram-2017). More specifically, influence functions (Garima-2020; Sui-2021; Kwon-2024) employ gradient-based analysis to help developers understand how models’ decisions are influenced by training data. Despite their clear benefits, a fundamental issue in influence computation is the computational overhead of utilizing the entire gradient space from all the layers of the LLM, which, for modern models, amount to billions of parameters.
@@
A common workaround (Figure 1) when employing influence functions in LLMs is to restrict the gradient input to only certain layers (Yen-2022; Chhabra-2025). However, there is little consensus on what the best layers are for influence estimation, and it is not known whether the output head, attention layers, or word embedding layers are better suited for this purpose. This is partly because very little work has been undertaken to analyze which layers are more useful for influence analysis. More specifically, one prior seminal work by Yen-2022 analyzes layer choice in computing language model data influence and finds that the first few (embedding) layers are most beneficial for influence estimation. However, as we discuss subsequently in our paper, these conclusions are derived from a strong assumption about the observed effectiveness of the layers being used, which might not hold in practice, requiring a deeper investigation into the effect of layers. Additionally, the work only studies one influence method and smaller-sized language models such as RoBERTa (Liu-2019-R), whereas we undertake our study across various LLM parameter sizes and recent models, thereby capturing general patterns of influence estimation as models scale.
@@
  • We evaluate the cancellation effect metric proposed by  Yen-2022 for identifying layers most suitable for influence estimation. Our theoretical analysis and experiments demonstrate that, when filtering large numbers of samples, the cancellation effect is an unreliable indicator.
  • We quantify the impact of detrimental training samples on language model behavior and identify the model layers most informative for influence estimation. We conduct extensive experiments using noisy versions of the GLUE benchmark as in prior influence work (Kwon-2024; Chhabra-2025) and employ several LLMs such as Llama-3.2, Qwen-2.5, and Mistral 7B.
@@ 2 Related Works
Our work falls within the broader area of interpretable machine learning, particularly focusing on gradient-method estimations of training data importance (Koh-2017; Yeh-2018; Yeh-2019; Jia-2019; Garima-2020; Khanna-2018; Sui-2021). In gradient-based models trained via empirical risk minimization, influence functions offer a practical means to estimate the impact of individual training samples without performing costly leave-one-out retraining. For deep learning, the foundational work (Koh-2017) introduced a Taylor-series approximation combined with the LiSSA optimization algorithm (Agarwal-2017) to make influence estimation tractable. Subsequent efforts, such as Representer Point (Yeh-2018) and Hydra (Chen-2021), aimed to improve these estimates, though their focus remained primarily on vision tasks. More recent advancements, including DataInf (Kwon-2024), Arnoldi iteration (Schioppa-2022), and Kronecker-factored curvature approximations (Grosse-2023), have scaled influence estimation to large generative language models like LLMs. Simpler methods have also emerged, using raw gradients as a proxy for influence  (Garima-2020; Charpiat-2019), often enhanced with ensemble-based strategies (Bae-2024; Kim-2023). Additionally, self-influence – computed solely within the training set – has been shown to be effective for identifying influential samples (Bejan-2023; Thakkar-2023). Influence functions have proven valuable across diverse applications, including classification (Koh-2017; Chhabra-2025; Chhabra-2024b), generative modeling (Kwon-2024; Schioppa-2022; Grosse-2023), active learning (Liu-2021b), in-context learning (Nguyen-2023; Van-2024; Askari-2025), and layer quality estimation (askari2025layerif).
@@
Gradient compensation in feature-attributed gradients has been studied previously (Liu-2020; Kapishnikov-2021). In data attribution, influence functions have been used to detect spurious training artifacts and improve performance (Han-2020; Pezeshkpour-2022; Askari-2025). Prior work (Hammoudeh-2024; Yen-2022) highlights that dot-product-based methods and high gradient cancellation may reduce the discriminative power of influence scores. We extend this line of work by empirically and theoretically investigating whether such cancellation impairs detection of mislabeled data.
@@ 3.1 Influence functions
The optimization step $\Delta\Theta_{1}$ in the direction of the vector given by the term under the parentheses decreases $f$ on $X^{\prime}$. The second term, $H^{-1}_{\Theta}\nabla l(\bar{x},\Theta)$ ($H$ - Hessian matrix) defines the optimization step $\Delta\Theta_{2}$ that would minimize the loss on the sample $\bar{x}$. Therefore, Eq. 1 estimates the alignment of these two optimization tasks.
The loss optimization on some training samples could be misaligned with the optimization of $f$. Therefore, fitting these samples would increase $f$. We refer to such samples as detrimental for the utility under consideration. This definition corresponds to the previously used definition in (Koh-2017; Kwon-2024; Chhabra-2025). In practice, it is not feasible to compute $I$ on large models due to the time complexity of the Hessian inversion (cubic time in the exact case). Therefore, influence is estimated approximately. The TracIn method (Garima-2020) replaces the Hessian vector product $\mathcal{H}=H^{-1}_{\Theta}\nabla l(\bar{x},\Theta)$ with the gradient $\nabla l(\bar{x},\Theta)$. Cosine similarity between the utility term and $\mathcal{H}\approx\nabla l(\bar{x},\Theta)$ serves as another first-order influence estimation approximation. DataInf (Kwon-2024) is a recently proposed second-order method based on swapping the order of matrix inversion and averaging across training samples in $\mathcal{H}$. Its error decreases when $|\Theta|$ is small, which makes this method suitable for fine-tuning with LoRA (Hu-2022; Dettmers-2023).
For cross-entropy loss, the Hessian is the block diagonal matrix of gradient second moments $|X|^{-1}\sum_{\bar{x}\in X}\nabla l(\bar{x},\Theta)\nabla l^{T}(\bar{x},\Theta)$, one block for one layer $l\in L$ in the neural network. Therefore, the influence $I$ is an aggregation of influence values across validation samples and model layers, defined as follows:
@@
To address the question of which layers are more effective at detecting detrimental samples, we also consider two influence functions proposed by Yen-2022. Both of them, $\operatorname{TracIn}_{we}$ and $\operatorname{TracIn}^{10}_{we}$, operate on word embeddings. The first, $\operatorname{TracIn}_{we}$, considers only the weights of the tokens (including special tokens) in both the training and the validation samples:
@@ 3.2 Cancellation effect
In their study of layer choice effects in influence estimation, Yen-2022 consider the compensation of gradients as contributed by training samples. They argue that if the sum of gradients w.r.t a particular weight $\theta$ becomes very small, then the optimizer would not change $\theta$ significantly because of them. However, the magnitudes of the summed training gradients could be very large and hence, the $\theta$-contributed influence score could also be large. The authors further state that $\theta$, the weight with a high cancellation effect, should be ignored in influence estimation because it reduces the model’s discrimination power. To this end, they propose a metric $C$ to estimate the cancellation effect on a layer $W$ and present the results where smaller $C$ values lead to better model confidence (Yen-2022):
@@ 4 Research questions
[RQ4]. How reliably can noise-distribution measures predict the downstream performance of influence-based data filtering methods? In Section 5.4, we study whether our proposed metrics: Noise Detection Rate (NDR) and Area Under the Curve (AUC), can serve as useful metrics for assessing the performance of influence functions, without requiring expensive fine-tuning.
@@ 5 Analysis and Experimental Results
Datasets. We evaluate detrimental sample filtering across eight GLUE datasets (Wang-2018): QNLI, MRPC, QQP, SST-2, CoLA, MNLI, RTE, and STS-B. These benchmarks cover a range of natural language understanding tasks, including sentence similarity, paraphrase detection, sentiment analysis, linguistic acceptability, and natural language inference. Together, they provide a diverse and widely used testbed for assessing model performance and the impact of data quality.
Models. The experiments are conducted on several LLM types: RoBERTa-Large (Liu-2019-R), LLaMA-3.2 1B (Touvron-2023), Qwen-2.5 1.5B (Qwen-25), and Mistral 7B (Albert-2023). These models represent various widely used transformer-based language models ranging from millions to several billions of parameters, offering a comprehensive view of influence behavior across model sizes and architectures.
@@
Roberta-Large
WE 2.2 $\pm$ 0.3 1 1 $\infty$ -0.3
@@
Methodology and Protocol. Similar to prior work (Kwon-2024; Chhabra-2025), we employ the five-stage pipeline detailed further in Appendix A): (1) Inject synthetic noise into training data and initialize from compressed embeddings. (2) Fine-tune on noisy data, selecting checkpoints with the lowest validation loss (Garima-2020), which we verify outperforms final-epoch or accuracy-based selection. (3) Compute influence values across all tunable layers. (4) Partition the model into embeddings (WE), attention groups (four splits), and classification head (CL), aggregating influence per training sample. (5) Remove 30% least influential samples and retrain, evaluating effectiveness by the best test accuracy. Each configuration is repeated over 10 seeds for robustness.
5.1 [RQ1] Verification of Assumptions about Cancellation Effect
@@
Table 1 presents the cancellation effect measured across LLMs, datasets, and seeds. The embeddings WE indeed have the smallest values. The median value of $C$ reflects the cancellation of an individual parameter, predominantly observed within a corresponding layer group. Value $C=1$ denotes the smallest cancellation when all the gradients are codirectional (e.g. majority of WE parameters). However, there are WE weights with infinite cancellation scores that have two or more collected gradients from training samples. Clearly, the results confirm that layer group cancellation $C$ “smoothes” the gradient compensation over a large number of parameters and hides those weights where the individual parameter gradient cancellation is high.
Yen-2022 showed that influence estimates from layers with lower $C$ better capture expected prediction changes for ground-truth classes when top-$k$ opponents are filtered out for a given sample. In contrast, we extend this analysis to the validation set as a whole, filtering opponents across many samples and measuring downstream performance via task accuracy. Results demonstrate only weak or no correlation between $C$ and performance (Table 1, column $\rho$: Spearman correlation).
Next, we provide theoretical evidence that reinforces the aforementioned issues with the cancellation effect and demonstrates how it is an unreliable estimator of influence performance.
@@
i.e. the separation between noisy and clean samples is strictly larger under $I_{\theta,\omega}$, thereby showing that contrary to the claim of Yen-2022, the inclusion of weights with high cancellation can improve influence estimates.
Proof. See Appendix G. ∎
@@
Next, owing to the limitations of the cancellation effect as previously discussed, we now seek to analyze which layers are better suited for influence estimation in language models. We compare layers based on how well they can detect detrimental samples in accordance with the framework outlined at the beginning of the section. In general, our results are only aligned with (Yen-2022) regarding the classification head CL performance. We observe that CL is highly susceptible to influential noise across models and methods. Dataset filtering with scores from CL improves downstream accuracy less than other layers (Figure 2).
@@
Figure 2: Best test accuracy of Mistral 7B after 30% filtering (averaged over 10 runs). Early and middle attention layers yield the strongest influence method performance on most tasks.
@@
Llama-3.2 1B. For Llama, influence functions perform worst among all settings: none of the configurations surpass the uniform removal baseline (Random), even when a large fraction of mislabeled samples is discarded. The strongest outcomes are obtained with DataInf on group 2 (layers 04–07, 0.64 win rate) and group 1 (layers 00–03, 0.58), comparable to Random (0.64). Accuracy trajectories (Figure 9, Appendix B) further show that Random outperforms influence methods on QQP, COLA, and STSB in later epochs. Despite this, noise distribution analysis (Figure 20, Appendix I) reveals non-uniform patterns, with mislabeled samples concentrating at both extremes of the influence range, describing a substantial fraction of detrimental samples as influential for the task goal.
@@
Remarks on RQ2 Findings. Across models and methods, we observe a consistent pattern in layer-level performance. With mean-score aggregation, the classification head (CL) shows limited ability to discriminate noise across all influence methods. DataInf and Cosine frequently outperform TracIn, particularly on early–middle attention layers of deeper models, suggesting that these layers encode representations that are most informative for noise filtering. Comprehensive rankings across all configurations are further provided in Appendix F (Tables 4, 5, 6).
@@
Next, we aim to study whether our proposed aggregation strategies (Section 3.3) can result in better influence function performance. We first analyze the correlation of training-sample influence scores across layers and methods (shown in Appendix E Figure 12), revealing three layer groups: early, middle, and late. Within a group, layers assign similar scores: for example, TracIn variations on WE consistently identify the same noisy samples, and late layers show strong agreement across all methods. Correlations between groups are weak, indicating that layers often compensate for one another. Across methods, DataInf closely aligns with TracIn on later layers, while Cosine shows low agreement, providing a complementary view of importance attribution. These patterns motivate the underlying ideas behind our proposed aggregation strategies, Rank and Vote, that reconcile layer- and method-level discrepancies. Figure 3 illustrates the performance differences relative to mean aggregation for Mistral 7B (results for other models show similar trends and are provided in Appendix H due to space limitations).
@@
Figure 4 further illustrates the proportion of noise filtered by each layer of Mistral 7B. Across models and influence functions, the Value B LoRA modules consistently capture the largest fraction of noisy samples, with peak filtering observed in the middle attention layers for larger models and later layers for smaller models (see Appendix E for additional results).

### v2

@@ Abstract
Identifying how training samples influence/impact Large Language Model (LLM) decision-making is essential for effectively interpreting model decisions and auditing large-scale datasets. Current training sample influence estimation methods (also known as influence functions) undertake this goal by utilizing information flow through the model via its first-order and higher-order gradient terms. However, owing to the large model sizes of today consisting of billions of parameters, these influence computations are often restricted to some subset of model layers to ensure computational feasibility. Prior seminal work by Yeh et al. (2022) in assessing which layers are best suited for computing language data influence concluded that the first (embedding) layers are the most informative for this purpose, using a hypothesis based on influence scores canceling out (i.e., the cancellation effect). In this work, we propose theoretical and empirical evidence demonstrating how the cancellation effect is unreliable, and that middle attention layers are better estimators for influence. Furthermore, we address the broader challenge of aggregating influence scores across layers, and showcase how alternatives to standard averaging (such as ranking and vote-based methods) can lead to significantly improved performance. Finally, we propose better methods for evaluating influence score efficacy in LLMs without undertaking model retraining, and propose a new metric known as the Noise Detection Rate (NDR) that exhibits strong predictive capability compared to the cancellation effect. Through extensive experiments across LLMs of varying types and scales, we concretely determine that the first (layers) are not necessarily better than the last (layers) for LLM influence estimation, contrasting with prior knowledge in the field.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated stellar performance on tasks across a number of applications and domains (Street et al., 2024; Mittelstädt et al., 2024; Marco et al., 2025). Despite these advancements, current state-of-the-art models still exhibit suboptimal behavior on complex reasoning tasks (Jiang et al., 2025), hallucinate responses and facts (Cleti and Jano, 2024), and can make biased and unfair decisions (Gallegos et al., 2024; Peters and Chin-Yee, 2025). To improve the trust and safety of LLMs, it is imperative to better interpret and understand model decision-making (Singh et al., 2024).
Recently, data attribution and influence methods (Hammoudeh and Lowd, 2024) have shown great promise in interpreting LLM behavior from the perspective of the training data (Grosse et al., 2023; Chhabra et al., 2025). These approaches seek to detect noisy, anomalous, out-of-distribution, or problematic training samples and conceptually link them to model output performance (Pleiss et al., 2020; Yang et al., 2024; Jiang et al., 2021). Even highly curated datasets can contain detrimental samples, often introduced unknowingly through human error or surreptitiously by malicious adversaries (Ekambaram et al., 2017). More specifically, influence functions (Pruthi et al., 2020; Sui et al., 2021; Kwon et al., 2024) employ gradient-based analysis to help developers understand how models’ decisions are influenced by training data. Despite their clear benefits, a fundamental issue in influence computation is the computational overhead of utilizing the entire gradient space from all the layers of the LLM, which, for modern models, amount to billions of parameters.
@@
A common workaround (Figure 1) when employing influence functions in LLMs is to restrict the gradient input to only certain layers (Yeh et al., 2022; Chhabra et al., 2025). However, there is little consensus on what the best layers are for influence estimation, and it is not known whether the output head, attention layers, or word embedding layers are better suited for this purpose. This is partly because very little work has been undertaken to analyze which layers are more useful for influence analysis. More specifically, one prior seminal work by Yeh et al. (2022) analyzes layer choice in computing language model data influence and finds that the first few (embedding) layers are most beneficial for influence estimation. However, as we discuss subsequently in our paper, these conclusions are derived from a strong assumption about the observed effectiveness of the layers being used, which might not hold in practice, requiring a deeper investigation into the effect of layers. Additionally, the work only studies one influence method and smaller-sized language models such as RoBERTa (Liu et al., 2019), whereas we undertake our study across various LLM parameter sizes and recent models, thereby capturing general patterns of influence estimation as models scale.
@@
  • We evaluate the cancellation effect metric proposed by  Yeh et al. (2022) for identifying layers most suitable for influence estimation. Our theoretical analysis and experiments demonstrate that, when filtering large numbers of samples, the cancellation effect is an unreliable indicator.
  • We quantify the impact of detrimental training samples on language model behavior and identify the model layers most informative for influence estimation. We conduct extensive experiments using noisy versions of the GLUE benchmark as in prior influence work (Kwon et al., 2024; Chhabra et al., 2025) and employ several LLMs such as Llama-3.2, Qwen-2.5, and Mistral 7B.
@@ 2 Related Works
Our work falls within the broader area of interpretable machine learning, particularly focusing on gradient-method estimations of training data importance (Koh and Liang, 2017; Yeh et al., 2018; 2019; Jia et al., 2019; Pruthi et al., 2020; Khanna et al., 2018; Sui et al., 2021). In gradient-based models trained via empirical risk minimization, influence functions offer a practical means to estimate the impact of individual training samples without performing costly leave-one-out retraining. For deep learning, the foundational work (Koh and Liang, 2017) introduced a Taylor-series approximation combined with the LiSSA optimization algorithm (Agarwal et al., 2017) to make influence estimation tractable. Subsequent efforts, such as Representer Point (Yeh et al., 2018) and Hydra (Chen et al., 2021), aimed to improve these estimates, though their focus remained primarily on vision tasks. More recent advancements, including DataInf (Kwon et al., 2024), Arnoldi iteration (Schioppa et al., 2022), and Kronecker-factored curvature approximations (Grosse et al., 2023), have scaled influence estimation to large generative language models like LLMs. Simpler methods have also emerged, using raw gradients as a proxy for influence  (Pruthi et al., 2020; Charpiat et al., 2019), often enhanced with ensemble-based strategies (Bae et al., 2024; Kim et al., 2023). Additionally, self-influence – computed solely within the training set – has been shown to be effective for identifying influential samples (Bejan et al., 2023; Thakkar et al., 2023). Influence functions have proven valuable across diverse applications, including classification (Koh and Liang, 2017; Chhabra et al., 2025; 2024), generative modeling (Kwon et al., 2024; Schioppa et al., 2022; Grosse et al., 2023), active learning (Liu et al., 2021), in-context learning (Nguyen and Wong, 2023; S. et al., 2024; Askari et al., 2025a), and layer quality estimation (Askari et al., 2025b).
@@
Gradient compensation in feature-attributed gradients has been studied previously (Liu et al., 2020; Kapishnikov et al., 2021). In data attribution, influence functions have been used to detect spurious training artifacts and improve performance (Han et al., 2020; Pezeshkpour et al., 2022; Askari et al., 2025a). Prior work (Hammoudeh and Lowd, 2024; Yeh et al., 2022) highlights that dot-product-based methods and high gradient cancellation may reduce the discriminative power of influence scores. We extend this line of work by empirically and theoretically investigating whether such cancellation impairs the detection of mislabeled data.
It is also interesting to note that prior work has studied the applicability of influence functions in LLMs, with varying results. Clearly, while influence functions have been used successfully across several LLM tasks and applications Wang et al. (2025); Zhang et al. (2025); Askari et al. (2025b); Xia et al. (2024), recent work Li et al. (2025) finds that Hessian-Free, DataInf, and LiSSA perform poorly on tasks such as harmful data detection, class attribution, and backdoor identification, due to iHVP approximation errors, uncertain convergence, and weak correlation between parameter changes and model behavior. Their experiments use default influence-function implementations with mean aggregation across all layers. In contrast, we extend these insights by identifying specific layers and aggregation strategies that achieve performance capable of exceeding the default settings in Li et al. (2025). As part of future work, we believe indetifying optimal layers for influence analysis can significantly improve the efficacy and fidelity of influence functions in LLMs, as observable in our main paper results.
Moreover, given our focus on where influence scores are strongest and how individual valuations can be effectively combined, we relate our analysis to knowledge-editing methods that similarly perform localization and aggregation. ROME (Meng et al., 2022) and R-ROME (Gupta et al., 2024a) use causal tracing to identify the MLP layer encoding a fact and apply a rank-one edit. MEMIT (Meng et al., 2023) generalizes this via per-layer causal-effect scores and top-$k$ distributed edits, while EMMET (Gupta et al., 2024b) refines aggregation to reduce redundancy. In contrast, we use gradient-based influence to locate layers most affected by beneficial or harmful training samples and, consistent with KE studies, find the strongest discriminative signals in middle layers.
@@ 3.1 Influence functions
The optimization step $\Delta\Theta_{1}$ in the direction of the vector given by the term under the parentheses decreases $f$ on $X^{\prime}$. The second term, $H^{-1}_{\Theta}\nabla l(\bar{x},\Theta)$ ($H$ - Hessian matrix) defines the optimization step $\Delta\Theta_{2}$ that would minimize the loss on the sample $\bar{x}$.
The loss optimization on some training samples could be misaligned with the optimization of $f$. Therefore, fitting these samples would increase $f$. We refer to such samples as detrimental for the utility under consideration. This definition corresponds to the previously used definition in (Koh and Liang, 2017; Kwon et al., 2024; Chhabra et al., 2025). In practice, it is not feasible to compute $I$ on large models due to the time complexity of the Hessian inversion (cubic time in the exact case). Therefore, influence is estimated approximately. The TracIn method (Pruthi et al., 2020) replaces the Hessian vector product $\mathcal{H}=H^{-1}_{\Theta}\nabla l(\bar{x},\Theta)$ with the gradient $\nabla l(\bar{x},\Theta)$. Cosine similarity between the utility term and $\mathcal{H}\approx\nabla l(\bar{x},\Theta)$ serves as another first-order influence estimation approximation. DataInf (Kwon et al., 2024) is a recently proposed second-order method based on swapping the order of matrix inversion and averaging across training samples in $\mathcal{H}$. Its error decreases when $|\Theta|$ is small, which makes this method suitable for fine-tuning with LoRA (Hu et al., 2022; Dettmers et al., 2023).
For cross-entropy loss, the Hessian is the block diagonal matrix of gradient second moments $|X|^{-1}\sum_{\bar{x}\in X}\nabla l(\bar{x},\Theta)\nabla l^{T}(\bar{x},\Theta)$, one block for one layer $l\in L$ in the neural network. Thus, $I$ is an aggregation of influence values across validation samples and model layers, defined as follows:
@@
To address the question of which layers are more effective at detecting detrimental samples, we also consider two influence functions proposed by Yeh et al. (2022). Both of them, $\operatorname{TracIn}_{we}$ and $\operatorname{TracIn}^{10}_{we}$, operate on word embeddings. The first, $\operatorname{TracIn}_{we}$, considers only the weights of the tokens (including special tokens) in both the training and the validation samples:
@@ 3.2 Cancellation effect
In their study of layer choice effects in influence estimation, Yeh et al. (2022) consider the compensation of gradients as contributed by training samples. They argue that if the sum of gradients w.r.t a particular weight $\theta$ becomes very small, then the optimizer would not change $\theta$ significantly because of them. However, the magnitudes of the summed training gradients could be very large and hence, the $\theta$-contributed influence score could also be large. The authors further state that $\theta$, the weight with a high cancellation effect, should be ignored in influence estimation because it reduces the model’s discrimination power. To this end, they propose a metric $C$ to estimate the cancellation effect on a layer $W$ and present the results where smaller $C$ values lead to better model confidence (Yeh et al., 2022):
@@ 4 Research questions
[RQ4]. How reliably can the detrimental sample distribution measures predict the downstream performance of influence-based data filtering methods? In Section 5.4, we study whether our novel proposed metrics: Noise Detection Rate (NDR) and Area Under the Curve (AUC), can serve as useful metrics for assessing the performance of influence functions, without requiring expensive fine-tuning.
@@ 5 Analysis and Experimental Results
Datasets. We evaluate detrimental sample filtering across eight GLUE datasets (Wang et al., 2018): QNLI, MRPC, QQP, SST-2, CoLA, MNLI, RTE, and STS-B. These benchmarks cover a range of natural language understanding tasks, including sentence similarity, paraphrase detection, sentiment analysis, linguistic acceptability, and natural language inference. Together, they provide a diverse and widely used testbed for assessing model performance and the impact of data quality.
Models. The experiments are conducted on several LLM types: RoBERTa-Large (Liu et al., 2019), LLaMA-3.2 1B (Touvron et al., 2023), Qwen-2.5 1.5B (Qwen et al., 2025), and Mistral 7B (Jiang et al., 2023). These models represent various widely used transformer-based language models ranging from millions to several billions of parameters, offering a comprehensive view of influence behavior across model sizes and architectures.
Methodology and Protocol. Similar to prior work (Kwon et al., 2024; Chhabra et al., 2025), we employ the five-stage pipeline detailed further in Appendix A): (1) Inject synthetic noise into training data and initialize from compressed embeddings. (2) Fine-tune on noisy data, selecting checkpoints with the lowest validation loss (Pruthi et al., 2020), which we verify outperforms final-epoch or accuracy-based selection. (3) Compute influence values across all tunable layers. (4) Partition the model into embeddings (WE), attention groups (four splits), and classification head (CL), aggregating influence per training sample. (5) Remove 30% least influential samples and retrain, evaluating effectiveness by the best test accuracy. Each configuration is repeated over 10 seeds for robustness.
5.1 [RQ1] Verification of Assumptions about Cancellation Effect
@@
Roberta-Large WE 2.2 $\pm$ 0.3 1 1 $\infty$ -0.3
@@
Table 1 presents the cancellation effect measured across LLMs, datasets, and seeds. The embeddings WE indeed have the smallest values. The median value of $C$ reflects the cancellation of an individual parameter, predominantly observed within a corresponding layer group. Value $C=1$ denotes the smallest cancellation when all the gradients are codirectional (e.g. majority of WE parameters). However, there are WE weights with infinite cancellation scores that have two or more collected gradients from training samples. Clearly, the results confirm that layer group cancellation $C$ “smoothes” the gradient compensation over a large number of parameters and hides those weights where the individual parameter gradient cancellation is high. Yeh et al. (2022) showed that influence estimates from layers with lower $C$ better capture expected prediction changes for ground-truth classes when top-$k$ opponents are filtered out for a given sample. In contrast, we extend this analysis to the validation set as a whole, filtering opponents across many samples and measuring downstream performance via task accuracy. Results demonstrate only weak or no correlation between $C$ and performance (Table 1, column $\rho$: Spearman correlation).
Next, we provide theoretical evidence that reinforces the aforementioned issues with the cancellation effect and demonstrate how it is an unreliable estimator of influence performance (see Appendix G for the proof).
@@
i.e. the separation between noisy and clean samples is strictly larger under $I_{\theta,\omega}$, showing that contrary to the claim of Yeh et al. (2022), the inclusion of high cancellation weights can improve influence.
@@
Next, owing to the limitations of the cancellation effect as previously discussed, we now seek to analyze which layers are better suited for influence estimation in language models. We compare layers based on how well they can detect detrimental samples in accordance with the framework outlined at the beginning of the section. In general, our results are only aligned with (Yeh et al., 2022) regarding the classification head CL performance. We observe that CL is highly susceptible to influential noise across models and methods. Dataset filtering with scores from CL improves downstream accuracy less than other layers (Figure 2).
@@
Figure 2: Best test accuracy of Mistral 7B after 30% filtering (averaged over 10 runs). Early and middle attention layers yield the strongest influence method performance on most tasks. Differences are statistically significant (Friedman p-value: 1e-142).
@@
Llama-3.2 1B. For Llama, influence functions perform worst among all settings: none of the configurations surpass the uniform removal baseline (Random), even when a large fraction of mislabeled samples is discarded. The strongest outcomes are obtained with DataInf on group 2 (layers 04–07, 0.64 win rate) and group 1 (layers 00–03, 0.58), comparable to Random (0.64). Accuracy trajectories (Figure 10, Appendix B) further show that Random outperforms influence methods on QQP, COLA, and STSB in later epochs. Despite this, noise distribution analysis (Figure 21, Appendix I) reveals non-uniform patterns, with mislabeled samples concentrating at both extremes of the influence range, describing a substantial fraction of detrimental samples as influential for the task goal. We provide additional discussion on Llama results in Appendix M.
@@
Remarks on RQ2 Findings. Across models and methods, we observe a consistent pattern in layer-level performance. With mean-score aggregation, the classification head (CL) shows limited ability to discriminate noise across all influence methods. DataInf and Cosine frequently outperform TracIn, particularly on early–middle attention layers of deeper models, suggesting that these layers encode representations that are most informative for noise filtering. Comprehensive rankings across all configurations are further provided in Appendix F (Tables 4, 5, 6). Additionally, we verify and reconfirm our findings on autoregressive datasets in Appendix J.
@@
Next, we aim to study whether our proposed aggregation strategies (Section 3.3) can result in better influence function performance. We first analyze the correlation of training-sample influence scores across layers and methods (shown in Appendix E Figure 13), revealing three layer groups: early, middle, and late. Within a group, layers assign similar scores: for example, TracIn variations on WE consistently identify the same noisy samples, and late layers show strong agreement across all methods. Correlations between groups are weak, indicating that layers often compensate for one another. Across methods, DataInf closely aligns with TracIn on later layers, while Cosine shows low agreement, providing a complementary view of importance attribution. These patterns motivate the underlying ideas behind our proposed aggregation strategies, Rank and Vote, that reconcile layer- and method-level discrepancies. Figure 3 illustrates the performance differences relative to mean aggregation for Mistral 7B (results for other models show similar trends and are provided in Appendix H).
@@
The positional voting scheme introduces the hyperparameter $k$ in Equation 10, representing the number of votes each validation sample assigns to training samples. We also study how aggregation performance varies as a function of $k$. Figure 4 shows the percentage of filtered detrimental samples across different values of $k$ and model layers for Qwen-2.5 1.5B with LoRA Value B module, which provides the strongest overall performance. For both DataInf and TracIn, performance is maximized for $k\in[10,50]$. In contrast, the Cosine-based method attains a local minimum around $k\approx 50$ and reaches its maximum at larger $k$. Similar trends are observed for other LoRA modules (see Appendix L for more details).
Figure 4: Filtered detrimental samples % with LoRA Value B module for Qwen-2.5 1.5B as a function of the number of votes $k$ (positional voting, Equation 10) and the selected model attention layer. The heatmaps on the axes show where performance is maximized.
@@
Figure 5 further illustrates the proportion of noise filtered by each layer of Mistral 7B. Across models and influence functions, the Value B LoRA modules consistently capture the largest fraction of noisy samples, with peak filtering observed in the middle attention layers for larger models and later layers for smaller models (see Appendix E for additional results). Additional sample-influence ranking variation analysis in Appendix K indicates that average noise influence attains its minimum at layers with the highest noise-detection efficiency.